Find an equation of the line containing the given pair of points
(1/4,-1/2) and (3/4,2)
What is the equation of the line?
y=?
Simplify your answer
Can anyone help please

Answer by Alan3354(69443)

Find an equation of the line containing the given pair of points
(1/4,-1/2) and (3/4,2)
What is the equation of the line?
Find the slope, m
m = diffy/diffx = (5/2)/(1/2)
m = 5
--------
Use y = mx + b with either point to find b, the y-intercept
2 = 5*(3/4) + b
b = -7/4
y = 5x - (7/4) slope-intercept form
20x - 4y = 7 standard form
--------------
Simplify your answer
Simplify in what way?